OpenStudy (anonymous):

write the equation with a slope of 3 through the piont(1,4) in simplest form. and how do i put it in simplest form?

OpenStudy (anonymous):

Use point-slope form. \[y – y_1 = m(x – x_1)\] It would become y\[y - 4 = 3(x - 1)\] which becomes \[y = 3x + 1\]